Test of the package rectopma and suggestions for its use (I need a long title)^{∗}

B.Benciolini^{†}  No Second Author^{‡}

1 Introduction

The package rectopma makes it possible to reuse the main content of \title and \author in different parts of a document. It is also possible to force linebreaks in the title with a command that is only active inside the top-matter, not when the title is re-printed elsewhere.

2 Instruction

The new commands \intitlebreak and \intitlebreakvs (vs= vertical skip) are used to force a line break in the title that disappears when the title itself is re-used outside the top matter of the paper. The content of \title and \author must be saved with the command \SaveTopMatter before the action of \maketitle and can be reprinted with the new commands \SavedAuthor and \SavedTitle. When they are reprinted the names of the various authors are simply separated by commas and the content of \thanks is ignored.

3 Examples

The authors of this paper are listed here after: B.Benciolini, No Second Author. This is obtained with \SavedAuthor. It is also possible to reprint the title, by means of \SavedTitle, and the result is: Test of the package rectopma and suggestions for its use (I need a long title). The title can be printed with a different style, as in: Test of the package rectopma and suggestions for its use (I need a long title), obtained with \textbf{\SavedTitle}.

∗Comments are welcome !

†e-mail: battista.benciolini@ing.uitn.it

‡No-Where Institute